# Bose–Fermi $N$-polaron state emergence from correlation-mediated blocking of phase separation

### v2

@@ Abstract
We study $N$ fermionic impurities in a one-dimensional lattice bosonic bath at unit filling. Using DMRG and mixed boundary conditions—closed for bosons, open for fermions— we find an N-polaron ground state replacing phase separation at high interspecies repulsion. This tightly bound state of clustered particles emerges due to strong impurity-bath correlations which induce large impurity-impurity correlations, that we quantify via the von Neumann entropy and bipartite mutual information respectively. This system also reveals a fermionic self-localization effect from a Mott insulator background due to local correlations between the impurities and the bath. The growth of long-range correlations breaks this Mott phase, resulting in the transition to localized impurity clusters. Finally, we show that there is a critical impurity number, which depends on intraspecies bosonic interaction, beyond which phase separation is recovered.
@@
Since the seminal work by Landau, impurities inside their surrounding many-body media can be understood as entities wearing a “dress” of background excitation, leading to a notion of quasi-particles known as polarons [1, 2, 3]. Studying the properties of polarons is often the first step to tackling more complicated many-body systems [4, 5], providing further insight that could lead to theoretical developments in condensed matter systems, such as the mechanism of the high-temperature superconductivity [6, 7, 8, 9, 10, 11], one of the ultimate goals of modern many-body physics. As exemplified by this system, the presence of strong correlations among particles may lead to various phenomena that are fundamentally and practically important. The realization of polarons in cold atoms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], opens a new and ideal platform to study correlation effects due to the high degree of tunability and clean environment [22, 23, 24, 25, 26, 27, 28, 29, 30]. This achievement has led to the study of the strong impurity-bath coupling regime, i.e. strongly correlated polarons, previously unattainable or uncontrollable in solid-state systems [31, 32, 33, 34, 35, 36, 37, 38, 39, 40], along with the investigation of the interplay between different particle statistics such as the Bose or Fermi polaron, an impurity immersed in a bosonic or fermionic bath [41, 42, 43, 44, 45], respectively, and multiple impurities where statistics and many-body effects play an important role [46, 47, 48, 49, 50, 51, 52].
@@
In this work, we study a one-dimensional lattice system that consists of a few fermionic impurities immersed in a bosonic bath at unit-filling. This system falls under the umbrella of Bose–Fermi mixtures, for which phase separation is known to occur for strong interspecies repulsion [53, 54, 55, 56] and which is also true for some few-impurity systems [57, 46, 58]. In the latter, boundary and finite size effects significantly affect the ground state of the full many-body system which is the focus of our work. We elucidate this by comparing open boundary conditions (BCs) for both bosonic and fermionic systems (Open(B)-Open(F) BCs) with closed boundary conditions for bosons while keeping the fermionic boundary open (Closed(B)-Open(F) BCs). While the miscible-immiscible phase transition is observed as expected for Open(B)-Open(F) BCs, for high population imbalance the Closed(B)-Open(F) BCs system avoids phase separation at large interspecies interactions and instead forms a $N$-polaron state bounded by large bath-impurity correlations. We characterize this state using density profiles and appropriate correlation functions, along with tools developed in quantum information theory such as the Von Neumann entropy and mutual information, which are used extensively in the characterization of quantum phase transitions [59, 60, 61, 62, 63].
@@
We consider a mixture of bosons and spin-polarized fermions loaded in a one-dimensional optical lattice of length $L$. The lattice depth is large enough that the low energy properties of the system can be well-described by a Bose–Fermi Hubbard model [64] with equal hopping parameters for the bosonic and fermionic species [65, 66]
@@
where $\hat{b}_{j}$ $(\hat{b}^{\dagger}_{j})$ and $\hat{f}_{j}$ $(\hat{f}^{\dagger}_{j})$ are the annihilation (creation) operators for bosons and fermions at the $j$-th site, respectively, with corresponding number operators $\hat{n}^{\rm B}_{j}=\hat{b}^{\dagger}_{j}\hat{b}_{j}$ and $\hat{n}^{\rm F}_{j}=\hat{f}^{\dagger}_{j}\hat{f}_{j}$. The terms $U_{\rm BB}$ and $U_{\rm BF}$ parameterize the strength of the boson-boson and boson-fermion coupling respectively, and will be varied to explore the phase diagram of the coupled system. In the following, we fix the number of bosons at unit-filling, $N_{\rm B}=L$, with system size $L=30$, while the number of fermions is much smaller than the background bosons, $N_{\rm F}\ll N_{\rm B}$. We initially fix $N_{\rm F}=2$ to illustrate the main physics and later discuss larger ratios of $N_{\rm F}/N_{\rm B}$. The ground state of the Hamiltonian (1) is found by using the Density Matrix Renormalization Group (DMRG) technique [67].
@@
Figure 1: Local observables and bosonic energy gaps. Left and right rows correspond to Open(B)-Open(F) and Closed(B)-Open(F) boundary conditions, respectively. Panel (a) and (d): Density profile of the bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ and fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ species as a function of $U_{\rm{BF}}/U_{\rm{BB}}$ at $U_{\rm{BB}}=6$. Panel (b) and (e) show the normalized overlap $\Lambda/\Lambda_{0}$, and (e) and (f) the bosonic gap $\Delta_{\rm B}$ in logarithmic scale as a function of $U_{\rm{BF}}$ and $U_{\rm{BB}}$. For better visibility, we show the density cuts for the fermionic (blue circles) and bosonic (red squares) profiles at $U_{\rm{BF}}/U_{\rm{BB}}=0$ and $2$ in the upper panel of (a) and (d). Green and blue lines indicate the coupling strengths that give the maximum of the interspecies entropy $S_{\rm BF}$ (see Fig. 3) and fermionic relative distance $D_{\rm FF}$ [67], respectively. The dotted black lines indicate $U_{\rm BB}=U_{\rm BF}$ for reference.
@@
We start by summarizing the case of Open(B)-Open(F) BCs and the well-known miscible-immiscible transition. In Fig. 1(a) we show the one-body density profiles of the mixture of the bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ and fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ species as a function of the interspecies coupling $U_{\rm BF}/U_{\rm BB}$ at fixed $U_{\rm BB}=6$. In the decoupled system ($U_{\rm{BF}}=0$, see left density cut from Fig. 1(a)), the background corresponds to a bosonic Mott insulator (MI) as the boson-boson interaction is much larger than the critical value of $U_{\rm{BB},0}^{\rm c}\approx 3.28$ [68, 69, 70, 71, 72] [^1] Although this critical point is measured in the thermodynamic limit, it is still an important reference value for the transition in this finite-sized system. , while the fermions are free and display two distinct peaks due to the Pauli repulsion. For sufficiently small interspecies interaction, $U_{\rm BF}\ll U_{\rm BB}$, the rigid insulating nature of the MI prevents any notable distortion to both density profiles; however, the density profiles change dramatically when $U_{\rm BF}\geq U_{\rm BB}$ as the system becomes phase separated with the fermions being individually localized at the box edges (see the right density cut in Fig. 1(a)).
@@
The nature of the transition alters dramatically if one considers Closed(B)-Open(F) BCs, as seen from the density profiles in Fig. 1(d). In particular, the most relevant change occurs for $U_{\rm BF}\gtrapprox U_{\rm BB}$ where the absence of a boundary effect for the bosonic cloud prevents the system from transitioning to an immiscible state. Instead, the fermions reorganize themselves inside the bulk of the lattice on top of a reduced but relatively large bosonic accumulation (see right density cut from Fig. 1(d)). The absence of the immiscible state can be captured by $\Lambda$ (see Fig. 1(e)), where for fixed $U_{\rm BB}$ and increasing $U_{\rm BF}$, the overlap decreases smoothly rather than the abrupt transition in the Open(B)-Open(F) BCs case and saturates at a finite and relatively large value. Indeed, the density of the fermions does not change with increasing strong boson-fermion coupling $U_{BF}/U_{BB}\ggg 1$ (we have checked up to $U_{BF}/U_{BB}=20$, see Supp. Matt. [67]) with the fermions remaining spread over many lattice sites and any further localization being suppressed.
As is the case for the miscible-immiscible transition, the SF bosonic background is necessary for this re-ordering to occur, as indicated in Fig. 1(f), where $\Delta_{\rm B}$ behaves similarly to the Open(B)-Open(F) BCs case since this is a bulk property. Here, we note that the MI to SF transition is signaled by the maximum of the relative distance between the two fermions $D_{\rm FF}=\left(N_{\rm F}^{2}-N_{\rm F}\right)^{-1}\sum_{j,k}^{L}|k-j|\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}^{\dagger}_{k}\hat{f}_{k}\hat{f}_{j}}$ [57, 67] and is indicated by the blue lines in Fig. 1(d-f), marking the MI-SF transition of the bosonic system in $\Delta_{B}$ and bounding the region of maximum miscibility in $\Lambda$. In the MI phase, bosonic correlations are short-ranged [76], and therefore bosons only couple locally to the fermions, leading to weak self-localization of individual impurities as can be seen from Fig. 1(d). After the transition, long-range correlations that characterize the SF bath can, in turn, mediate long-range interactions between the impurities which leads them to localize together around the center of the lattice [67].
@@
$\hat{\rho}_{\rm F}^{(2)}$ $=$ $\frac{1}{N_{\rm F}(N_{\rm F}-1)}\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}^{\dagger}_{m}\hat{f}_{l}\hat{f}_{k}},$ (8)
respectively. The one-body reduced density matrix is an auto-correlation function describing the probability of finding the same fermion simultaneously in two different lattice sites $j$ and $k$, while the two-body reduced density matrix describes the joint probability of finding one fermion in sites $j$ and $k$ and the other fermion in sites $l$ and $m$. As the latter is difficult to visualize we instead show in Fig. 2(a-c) its diagonal contribution, i.e. when $j=m$ and $k=l$, via the two-body density profile $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle=(1-\delta_{j,k})\langle\hat{n}^{\rm F}_{j}\rangle\langle\hat{n}^{\rm F}_{k}\rangle$. This represents the probability of finding one fermion at site $j$ and the other at site $k$, and is, of course, zero at the diagonal ($j=k$) due to the Pauli exclusion principle. Free fermions are anti-correlated, as shown in Fig. 2(a), and the two particles are separated into two different sides of the box potential.
Figure 2: (a-c) Two-body fermionic density profiles $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle$ and (d-f) one-body fermionic reduced density matrix $\hat{\rho}_{\rm F}^{(1)}$ for Closed(B)-Open(F) BCs. Here we have fixed the boson-boson coupling $U_{\rm BB}=6$ and take (a,d) $U_{\rm BF}=0$, (b,e) $U_{\rm BF}=1.333\,U_{\rm BB}$, and (c,f) $U_{\rm BF}=2\,U_{\rm BB}$.
However, for finite interspecies coupling $U_{\rm BF}=1.333U_{\rm BB}$ (Fig. 2(b)), for which the bosons are in the SF phase, the fermions start to bunch (they are more likely to be found near one another) and the relative distance between them decreases. For even larger $U_{\rm BF}$ (Fig. 2(c)) the fermions form a tightly bound state and $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle$ is significantly broadened along the $j=k$ direction, indicating a large variance in the “center of mass” position and enhanced particle bunching (the fermions are most likely to be found in adjacent lattice sites). A similar behaviour can be seen in the one-body density matrix shown in Fig. 2(d-f), with the tightly bound state only displaying short-range off-diagonal coherences in Fig. 2(f). The emergence of this tightly bound fermionic state is due to effective attractive interactions that are mediated by the coupling to the bosonic background and are indicative of polaron physics [47, 50], showing that a bipolaron state is formed when considering Closed(B)-Open(F) BCs.
@@
Figure 3: (a,c) Interspecies entropy $S_{\rm BF}$ and (b,d) fermionic mutual information $I_{\rm FF}$ as a function of $U_{\rm BF}/U_{\rm BB}$ for fixed $U_{\rm BB}=2$ (blue dot-dashed), $4$ (purple dashed) and $6$ (orange solid), and $I_{\rm FF}$ is normalized to the non-interacting value $I_{\rm{FF},0}=I_{\rm FF}(U_{\rm BF}=0)$. (a,b) correspond to Open(B)-Open(F) and (c-f) to Closed(B)-Open(F) BCs. The inset in (d) shows $U_{\rm BF}<U_{\rm BB}$ for $U_{\rm BB}=6$ with highlighted MI (gray) and SF (white) background phases, separated by the vertical solid blue line from Fig. 1(d). (e) $I_{\rm FF}$ vs $N_{\rm F}$ with $U_{\rm BF}\to\infty$ and (f) $I_{\rm FF}$ vs $U_{\rm BF}/U_{\rm BB}$ for $N_{\rm F}=15$ and the same $U_{\rm BB}$ values and labels as (a-d). The inset in (f) exhibits the fermionic (blue dots) and bosonic (red squares) density profiles for $U_{\rm BB}=6$ and $U_{\rm BF}=24$. Gray dotted lines indicate $U_{\rm BF}=U_{\rm BB}$ (vertical) and $I_{\rm FF}=I_{\rm{FF},0}$ (horizontal) for reference.

### v3

@@ Abstract
We study $N$ fermionic impurities in a one-dimensional lattice bosonic bath at unit filling. Using DMRG and mixed boundary conditions—periodic for bosons, open for fermions— we find an $N$-polaron ground state replacing phase separation at high interspecies repulsion. This tightly bound state of clustered particles emerges due to strong impurity-bath correlations, which induce large impurity-impurity correlations that we quantify via the von Neumann entropy and bipartite mutual information, respectively. This system also reveals a fermionic self-localization effect from a Mott insulator background due to local correlations between the impurities and the bath. The growth of long-range correlations breaks this Mott phase, resulting in the transition to impurity clusters delocalized along the system. Finally, we show that there is a critical impurity number, which depends on intraspecies bosonic interaction, beyond which phase separation is recovered.
I Introduction
@@
Since the seminal work by Landau, impurities inside their surrounding many-body media can be understood as entities wearing a “dress” of background excitation, leading to a notion of quasi-particles known as polarons [1, 2, 3]. Studying the properties of polarons is often the first step to tackling more complicated many-body systems [4, 5], providing further insight that could lead to theoretical developments in condensed matter systems, such as the mechanism of high-temperature superconductivity [6, 7, 8, 9, 10, 11], one of the ultimate goals of modern many-body physics. As exemplified by this system, the presence of strong correlations among particles may lead to various phenomena that are fundamentally and practically important.
The realization of polarons in cold atoms [12, 13, 14, 15, 16, 17], opens a new and ideal platform to study correlation effects due to the high degree of tunability and clean environment [18, 19, 20, 21]. This achievement has led to the study of the strong impurity-bath coupling regime, i.e. strongly correlated polarons, previously unattainable or uncontrollable in solid-state systems [22, 23, 24], along with the investigation of the interplay between different particle statistics such as the Bose or Fermi polaron, an impurity immersed in a bosonic or fermionic bath respectively [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], and multiple impurities where statistics and many-body effects play an important role [46, 47, 48, 49, 50, 51, 52, 53, 54].
@@
In this work, we study a one-dimensional lattice system that consists of a few fermionic impurities immersed in a bosonic bath at unit-filling. This system falls under the umbrella of Bose–Fermi mixtures, for which phase separation is known to occur for strong interspecies repulsion [55, 56, 57, 58] and which is also true for some few-impurity systems [59, 46, 60]. In the latter, boundary conditions (BCs) and finite size effects significantly affect the ground state of the full many-body system which is the focus of our work. We elucidate this by comparing open BCs for both bosonic and fermionic systems (Open(B)-Open(F) BCs) with periodic BCs for bosons while keeping the fermionic BCs open (Periodic(B)-Open(F) BCs). While the miscible-immiscible phase transition is observed as expected for Open(B)-Open(F) BCs, for high population imbalance the Periodic(B)-Open(F) BCs system avoids phase separation at large interspecies interactions and instead forms a $N$-polaron state bounded by large bath-impurity correlations. We characterize this state using density profiles and appropriate correlation functions, along with tools developed in quantum information theory such as the von Neumann entropy and mutual information, which are used extensively in the characterization of quantum phase transitions [61, 62, 63, 64, 65].
Our work is organized as follows: First, we explain the details of the Bose-Fermi model (Sec. II), after which we illustrate the formation of the polaron and the differences that appear from mixed boundary conditions (Sec. III). Then, we discuss the relationship between the formation of polarons with the superfluid-insulator transition of the Bose Hubbard model (Sec. IV). We then elucidate the polaron structure at large interactions and characterize its formation using quantum information techniques (Sec. V). Finally, we discuss the role of population imbalance (Sec. VI) and system size (Sec. VII) on the mixture, and provide concluding remarks (Sec. VIII).
II Model
@@
We consider a mixture of bosons and spin-polarized fermions loaded in a one-dimensional optical lattice of length $L$. The lattice depth is large enough that the low energy properties of the system can be well-described by a Bose–Fermi Hubbard model [66]
@@
where $\hat{b}_{j}$ $(\hat{b}^{\dagger}_{j})$ and $\hat{f}_{j}$ $(\hat{f}^{\dagger}_{j})$ are the annihilation (creation) operators for bosons and fermions at the $j$-th site, respectively, with corresponding number operators $\hat{n}^{\rm B}_{j}=\hat{b}^{\dagger}_{j}\hat{b}_{j}$ and $\hat{n}^{\rm F}_{j}=\hat{f}^{\dagger}_{j}\hat{f}_{j}$. The hopping integral $t$ is assumed equal for both species, which has been realized in ${}^{171}\rm{Yb}-^{174}\rm{Yb}$ [67] and ${}^{6}\rm{Li}-^{7}\rm{Li}$ [68] experiments. The terms $U_{\rm BB}$ and $U_{\rm BF}$ parameterize the strength of the boson-boson and boson-fermion coupling respectively, tunable by the use of appropriate Feshbach resonances [69, 70, 71]. The BCs of each species are encoded in the hopping terms of their respective Hamiltonians (2) and (3): for periodic BCs an additional hopping term between the last and first site of the chain is added while for open BCs that term is omitted. This type of mixed BCs can be engineered in experiments using a ring trap to enforce the periodic BCs with a potential barrier [72, 73, 74] that breaks translational symmetry (open BCs) for only one species in the mixture employing species-dependent external potentials such as tuning-detuning schemes used for alkali atoms [75].
@@
In the following, we fix the number of bosons at unit-filling, $N_{\rm B}=L$, with system size $L=30$, while the number of fermions is much smaller than the background bosons, $N_{\rm F}\ll N_{\rm B}$, for which we fix $N_{\rm F}=2$. Later we consider larger ratios of $N_{\rm F}/N_{\rm B}$ and system sizes. The interaction strengths will be varied keeping $t=1$ setting the energy scale. To explore the phase diagram of the coupled system (1) the ground state is constructed using the Density Matrix Renormalization Group (DMRG) technique in the tensor network formalism [76]. Details of our numerical simulations are given in Appendix A.
III Polaron formation
We start by summarizing the case of Open(B)-Open(F) BCs and the well-known miscible-immiscible transition. In Fig. 1(a) we show the one-body density profiles of the mixture of the bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ and fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ species as a function of the interspecies coupling $U_{\rm BF}/U_{\rm BB}$ at fixed $U_{\rm BB}=6$.
@@ III Polaron formation
Figure 1: Local observables and bosonic energy gaps. The left and right rows correspond to Open(B)-Open(F) and Periodic(B)-Open(F) BCs, respectively. Panel (a) and (d): Density profile of the bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ and fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ species as a function of $U_{\rm{BF}}/U_{\rm{BB}}$ at $U_{\rm{BB}}=6$. Panel (b) and (e) show the normalized overlap $\Lambda/\Lambda_{0}$, and (c) and (f) the bosonic gap $\Delta_{\rm B}$ in logarithmic scale as a function of $U_{\rm{BF}}$ and $U_{\rm{BB}}$. Blue and green lines indicate the coupling strengths that give the maximum of the fermionic relative distance $D_{\rm FF}$ (see Fig. 3) and interspecies entropy $S_{\rm BF}$ (see Fig. 5,6), respectively. The dotted black lines indicate $U_{\rm BB}=U_{\rm BF}$ for reference.
@@
In the decoupled system ($U_{\rm{BF}}=0$) exhibited in Fig. 2(a), the background corresponds to a bosonic Mott insulator (MI) as the boson-boson interaction is much larger than the critical value of $U_{\rm{BB},0}^{\rm c}\approx 3.28$ [77, 78, 79, 80, 81]  [^1] Although this critical point is measured in the thermodynamic limit, it is still an important reference value for the transition in this finite-sized system., while the fermions are free and display two distinct peaks due to the Pauli repulsion.
Figure 2: Density profile of the fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ (blue circles) and bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ (red diamonds) species at $U_{\rm{BB}}=6$ with (a) Open(B)-Open(F) (b) and Periodic(B)-Open(F) BCs for $U_{\rm{BF}}=0$ (left column) and $U_{\rm{BF}}/U_{\rm{BB}}=2$ (right column).
For sufficiently small interspecies interaction, $U_{\rm BF}\ll U_{\rm BB}$, the rigid insulating nature of the MI prevents any notable distortion to both density profiles; however, the density profiles change dramatically when $U_{\rm BF}\geq U_{\rm BB}$ (see Fig. 1(a)) as the system becomes phase separated with the fermions being individually localized at the box edges, which we show explicitly for $U_{\rm{BF}}/U_{\rm BB}=2$ in Fig. 2(a).
@@
The nature of the transition alters dramatically if one considers Periodic(B)-Open(F) BCs, as seen from the density profiles in Fig. 1(d). In particular, the most relevant change occurs for $U_{\rm BF}\gtrapprox U_{\rm BB}$ where the absence of a boundary effect for the bosonic cloud (compare edges of the bosonic profile between BCs in Fig. 2 at $U_{\rm BF}=0$) prevents the system from transitioning to an immiscible state. Instead, the fermions cluster together and reorganize themselves inside the bulk of the lattice on top of a reduced but relatively large bosonic accumulation shown in Fig. 2(b) for $U_{\rm BF}/U_{\rm BB}=2$.
The absence of the phase separated state can be captured by $\Lambda$ (see Fig. 1(e)), where for fixed $U_{\rm BB}$ and increasing $U_{\rm BF}$, the overlap decreases smoothly rather than the abrupt transition in the Open(B)-Open(F) BCs case, and saturates at a finite and relatively large value. Indeed, the density of the fermions does not change with increasingly large boson-fermion coupling $U_{BF}/U_{BB}\gg 1$, with the fermions remaining spread over many lattice sites and any further localization being suppressed. In this regime the fermions have formed a tightly bound bipolaron state due to the strong mediated attractive interactions from the boson bath, which we will further characterize in the following sections.
Finally, as in the case of the miscible-immiscible transition, the SF bosonic background is necessary for this re-ordering of the fermions to occur, as indicated in Fig. 1(f), where $\Delta_{\rm B}$ behaves similarly to the Open(B)-Open(F) BCs case (Fig. 1(c)) since this is a bulk property.
IV SF-MI transition and fermionic relative distance
Besides the polaron formation, the Periodic(B)-Open(F) BCs exhibit an intrinsic relation between the SF-MI transition of the bosonic bath and the distribution of the fermionic impurities. This can be succinctly seen in the relative distance between the two fermions [59]
$D_{\rm FF}=\frac{1}{N_{\rm F}(N_{\rm F}-1)}\sum_{j,k}^{L}|k-j|\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}^{\dagger}_{k}\hat{f}_{k}\hat{f}_{j}}\,,$ (7)
shown in Fig. 3 as a function of $U_{\rm BF}/U_{\rm BB}$ for different $U_{\rm BB}$ values and BCs.
Figure 3: Fermionic relative distance $D_{\rm FF}$ as a function of $U_{\rm BF}/U_{\rm BB}$ for different $U_{\rm BB}$ values and with (a) Open(B)-Open(F) and (b) Periodic(B)-Open(F) BCs. The dot on each curve in panel (b) indicates the maximum distance for corresponding $U_{\rm BB}$. The dotted line references $U_{\rm BF}=U_{\rm BB}$.
As a reference, we first analyze the case for Open(B)-Open(F) BCs (Fig. 3(a)). Here the fermions appear to spread along the lattice progressively as we increase $U_{\rm BF}$ which agrees with the behavior at the miscible-immiscible transition ($U_{\rm BF}\approx U_{\rm BB}$) where the relative distance rapidly increases indicating the separation of the impurities. This increase is more abrupt for higher $U_{\rm BB}$ due to the breaking of the MI phase via the interaction with the fermions.
Periodic(B)-Open(F) BCs (Fig. 3(b)) present a different behavior for $U_{\rm BB}$ values below and above $U_{\rm BB,0}^{\rm c}$. In the former ($U_{\rm BB}=2$ in Fig. 3(b)), $D_{\rm FF}$ decreases as the fermions grow closer with increasing $U_{\rm BF}$ as a prelude of the polaron formation. On the other hand, for $U_{\rm BB}>U_{\rm BB,0}^{\rm c}$ ($U_{\rm BB}=6$, $10$ in Fig. 3(b)) before the polaron emergence at $U_{\rm BF}/U_{\rm BB}\approx 1$ there is an increase in $D_{\rm FF}$. Here as the background is in the MI phase, bosonic correlations are short-ranged [85], and therefore bosons only couple locally to the fermions, leading to weak self-localization of individual impurities that can be seen as an increase in the fermionic density maxima and the emergence of minima in the bosonic density in Fig. 1(d). After the SF-MI transition, long-range correlations that characterize the SF bath can, in turn, mediate interactions between the impurities longer than the MI bath which leads them to cluster together around the center of the lattice, decreasing $D_{\rm FF}$ as it is for $U_{\rm BB}<U_{\rm BB,0}^{\rm c}$. It is important to note that the fermionic self-localization effect occurs for both BCs, given that this is a phenomenon of the MI bulk, and is also seen in Fig. 1(a) for $U_{\rm BF}/U_{\rm BB}<1$.
Therefore, a characteristic maximum in $D_{\rm FF}$ appears when the impurities drive the SF-MI transition of the bosonic background’s phase, which is marked by a dot on each curve in Fig. 3(b). We highlight this maximum relative distance for each $U_{\rm BF}$ and $U_{\rm BB}$ with solid blue lines in Fig. 1(d-f). Remarkably, this clearly delineates the SF-MI phase boundary by agreeing with the other phase transition indicators, like the breaking of the bosonic uniform density profile $\expectationvalue{\hat{n}_{j}^{\rm B}}$ in Fig. 1(d), the decrease from maximum miscibility in the overlap $\Lambda$ in Fig. 1(e) and the vanishing of the energy gap $\Delta_{B}$ in Fig. 1(f). Moreover, we observe that by increasing $U_{\rm BB}$ the maximum of the relative distance in Fig. 3(b) also increases. This occurs because $\Delta_{\rm B}$ grows with $U_{\rm BB}$ as shown in Fig. 1(f), hence the MI phase becomes more stable against the perturbation by the impurities, which in turn, allows the impurities to self-localize further without breaking the MI state. We provide further characterization of the bosonic correlations in the transition in Appendix B.
V Strong coupling regime
@@
$\hat{\rho}_{\rm F}^{(2)}$ $=$ $\frac{1}{N_{\rm F}(N_{\rm F}-1)}\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}^{\dagger}_{k}\hat{f}_{m}\hat{f}_{l}},$ (9)
respectively. The one-body reduced density matrix is an auto-correlation function describing the probability of finding the same fermion simultaneously in two different lattice sites $j$ and $k$, and we have already studied the contribution of its diagonal ($j=k$) in the form of fermionic density profiles $\expectationvalue{\hat{n}^{\rm F}_{j}}$ in Fig. 1. On the other hand, the two-body reduced density matrix describes the joint probability of finding one fermion in sites $j$ and $l$ and the other fermion in sites $k$ and $m$, which was used to define $D_{\rm FF}$ in (7). As $\hat{\rho}_{\rm F}^{(2)}$ is difficult to visualize we instead study its diagonal contribution i.e. when $l=j$ and $m=k$, via the two-body density profile
$\expectationvalue{\hat{n}^{\rm FF}_{j,k}}$ $=\expectationvalue{\hat{f}^{\dagger}_{j}\hat{f}^{\dagger}_{k}\hat{f}_{k}\hat{f}_{j}}=\expectationvalue{\hat{f}^{\dagger}_{j}(\hat{f}_{j}\hat{f}^{\dagger}_{k}-\delta_{j,k})\hat{f}_{k}}$
$=(1-\delta_{j,k})\expectationvalue{\hat{n}^{\rm F}_{j}\hat{n}^{\rm F}_{k}},$ (10)
where we used the anticommutation properties of fermionic operators along with $\hat{n}^{\rm F}_{j}(\hat{n}^{\rm F}_{j}-1)=0$. This represents the probability of finding one fermion at site $j$ and the other at site $k$, and is, of course, zero at the diagonal ($j=k$) due to the Pauli exclusion principle.
We start with an analysis of the two-body density profile shown in Fig. 4(a-c) for $U_{\rm BB}=6$. Free fermions are anti-correlated (Fig. 4(a) for $U_{\rm BF}=0$) as the two particles are separated into two different sides of the box potential.
Figure 4: (a-c) Two-body fermionic density profiles $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle$ and (d-f) one-body fermionic reduced density matrix $\hat{\rho}_{\rm F}^{(1)}$ for Periodic(B)-Open(F) BCs. Here we have fixed the boson-boson coupling $U_{\rm BB}=6$ and take (a,d) $U_{\rm BF}=0$, (b,e) $U_{\rm BF}=1.333\,U_{\rm BB}$, and (c,f) $U_{\rm BF}=2\,U_{\rm BB}$.
However, for finite interspecies coupling $U_{\rm BF}=1.333U_{\rm BB}$ (Fig. 4(b)), which is after the MI-SF transition for the bosons and they are therefore in the SF phase, the fermions start to bunch (they are more likely to be found near one another) and the relative distance between them decreases (see also Fig. 3(b)). For even larger $U_{\rm BF}$ (Fig. 4(c)) the fermions form a tightly bound state and $\langle\hat{n}^{\rm{FF}}_{j,k}\rangle$ is significantly broadened along the $j=k$ direction, indicating a large variance in the center of mass position (they are spread across the lattice) and enhanced particle bunching (the fermions are most likely to be found in adjacent lattice sites). A similar behaviour can be seen in the one-body density matrix shown in Fig. 4(d-f), with the tightly bound state only displaying short-range off-diagonal coherences in Fig. 4(f). The emergence of this tightly bound fermionic state is due to effective attractive interactions that are mediated by the coupling to the bosonic background and are indicative of polaron physics [48, 51], confirming that a bipolaron state is formed when considering Periodic(B)-Open(F) BCs.
@@
Figure 5: (a) Interspecies entropy $S_{\rm BF}$ and (b) fermionic mutual information $I_{\rm FF}$ as a function of $U_{\rm BF}/U_{\rm BB}$ for $N_{\rm F}=2$, fixed $U_{\rm BB}=2$ (blue dot-dashed), $4$ (purple dashed) and $6$ (orange solid) and Open(B)-Open(F) BCs. $I_{\rm FF}$ is normalized to the non-interacting value $I_{\rm{FF},0}=I_{\rm FF}(U_{\rm BF}=0)$. Gray dotted lines indicate $U_{\rm BF}=U_{\rm BB}$ (vertical) and $I_{\rm FF}=I_{\rm{FF},0}$ (horizontal) for reference.
Figure 6: (a) Interspecies entropy $S_{\rm BF}$ and (b) fermionic mutual information $I_{\rm FF}$ as a function of $U_{\rm BF}/U_{\rm BB}$ for $N_{\rm F}=2$, fixed $U_{\rm BB}=2$ (blue dot-dashed), $4$ (purple dashed) and $6$ (orange solid) and Periodic(B)-Open(F) BCs. Inset in (b): Closeup of the mutual information across the SF-MI transition with the blue line obtained from Fig. 3. $I_{\rm FF}$ is normalized to the non-interacting value $I_{\rm{FF},0}=I_{\rm FF}(U_{\rm BF}=0)$. Gray dotted lines indicate $U_{\rm BF}=U_{\rm BB}$ (vertical) and $I_{\rm FF}=I_{\rm{FF},0}$ (horizontal) for reference.